The diagonal of a rectangle is 5cm & one of its side is 4 cm. Its area is?
Answer by mananth(16949)

diagonal = 5
one side = 4
let other side be x
..
they form a right triangle.
so use pythagoras theorem
diagonal is hypotenuse
..
x^2+4^2= 5^2
x^2+16=25
x^2= 25-16
x^2= 9
x= 3 the other sides
...
m.ananth@hotmail.ca